OpenStudy (anonymous):

Which is the exact area of a circle with radius fourteen centimeters? fourteen centimeters squared fourteen times pi centimeters squared twenty-eight times pi centimeters squared one hundred ninety-six times pi centimeters squared help

OpenStudy (goformit100):

Use the formula

OpenStudy (anonymous):

is it 196?

OpenStudy (mathstudent55):

\(A = \pi r^2 \) r = radius Just substitute in the formula and evaluate. Keep \(\pi\) as it is.

OpenStudy (mathstudent55):

\(A = 196\pi~cm^2 \)

OpenStudy (anonymous):

Which is the cost per square inch of a pizza with diameter 20 inches if the cost of the pizza is $16? Be sure to round your answer to the nearest cent. about 1 cent about 5 cents about twenty-five cents about eighty cents

OpenStudy (anonymous):

im confused ^

OpenStudy (mathstudent55):

@taliyahparker It is the answer that has 196 in it, but remember that the answer is \( 196\pi ~cm^2 \), not just simply 196.

OpenStudy (anonymous):

I got that @mathstudent55

OpenStudy (mathstudent55):

The area of a circle is calculated by multiplying the square of the radius by pi. Ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Here, the radius is 14 cm. The square of 14 cm = (14 cm)^2 = 196 cm^2. Then you multiply that by \(\pi\), so you get \(196\pi ~cm^2\). Do you understand now?

OpenStudy (anonymous):

yes but I was talking about that other problem when I said I was confused

OpenStudy (mathstudent55):

Oh, sorry, I had not noticed the second problem. I'll look at it now.

OpenStudy (mathstudent55):

You need to do 2 things. 1. First, find the area of the pizza in square inches. 2. Then convert the price of the pizza to cents and divide by the area of the pizza.

OpenStudy (mathstudent55):

Remember, the area of a circle is \(A = \pi r^2 \) Here we are given that the diameter of the pizza is 20 in. What is the radius of the pizza if the diameter is 20 in.?